• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Carlos

    Log entry time 23:17:42 on November 6,2004

    Entry number 136016

    keyword=Compton rates

    Compton rates were very high when the beam was back, so high that the central crystal PMT saturated and gave actually zero counts(!). Called Sirish and MCC.

    We lowered the central crystal HV to get rates values, and saw around 25kHz/uA, instead of the usual 1kHz/uA. MCC tried to optimise vertical and horizontal positions but got no improvement. Then Sirish suggested optimising pathlengths. No significant results. MMC called the expert on call in and he started looking into it. He finally managed to tune the beam to get the usual rates.

    We have now started production runs with the Compton.